# Economic thermodynamics and inflation

### v1

@@ I Introduction
In this study, it is shown that Raychaudhuri equation R32 ; R33 ; R34 ; R35 can describe an economic system without going into deep mathematical proofs. The solutions of the equation for adiabatic and non-adiabatic processes that occur in the energy conservation equation, which are considered in the thermodynamic framework, are investigated. At this point, it has been observed that the adiabatic process in which hyperinflation emerged can be eliminated by certain economic processes and monetary policies directing the system. In this direction, we consider the economic system as a field that can be defined as a scalar field $\phi$. The existence of a minimum inflation in this field is a necessity even at the beginning. The Raychaudhuri equation appears as an equation describing such an economic system that can bend and expand continuously. When we say bending or curvature of the field, we mean whether the unit (point A in Figure 1) with kinetic energy (monetary purchasing power) can reach a point with potential energy (movable-real estate, valuables) within the field.
@@ II The Raychaudhuri equation and inflation
In general, inflation is a concept used to explain the increase in average prices in the economy. On the other hand, this also means that money loses its value (Reserve Bank of New Zealand R36 ). In the mathematical constructs of economics, rather than theoretical proof, economic conceptual phenomena are structured within the framework of theory. With these intuitive and reasonable assumptions, economic systems are followed and analyzed through the lens of theory. We will mathematically restructure inflation dynamics, which expresses the constant increase in the price of movable and immovable goodsR31 , not in a part of the economic system but in its general structure, with a scalar field. Similar to cosmological inflation, we assume that scalar fields have a role in causing inflation in economy. In this context, we refer to ”developmental symmetry” as the property of a phenomenon in the observable universe having a structure that is similar to or symmetrical with a phenomenon in the social structure. Considering the fact that inflation takes place in the entire region of a real economics system, within the framework of thermodynamic boundary we assume that the inflation occurs and in this way is distributed throughout all point of the homogeneous structure (economics system).
Let the purchasing power be represented by point $A$ in the field and the price of the good by point $B$. The distance between both points is scaled by $x$. However, we consider the condition of availability of $B$ in an economic system (no deflation). It is worth emphasizing that when the acceleration term is $\ddot{x}>0$ it states the field is in the inflation phase and at $\ddot{x}<0$ it is in the deflation phase. Here, it should be apprehensible that an increase in the x scale under the $\ddot{x}>0$ condition indicates a decrease in purchasing power or shopping, thus indicating inflation. In the literature, the relative rate of change in the price of the good is given by the equality $C=\frac{\dot{p}}{p}$R37 ; R38 ; R39 ; R40 . According to the method we have discussed in this study, we give another definition of the inflation rate of change. For this purpose, one can define inflation rate $C=\frac{\dot{x}}{x}$ that shows how an economic system or inflation field has evolved. In such a system, the emergence of a developed economic equilibrium depends on the system’s conversion of $B$, which is formed by production, as added value to the purchasing power at point $A$. In an inflationary system, the most important way of converting movable-immovable goods (including technological production) into monetary resources is exportation. Another known tool is the resources transferred to the system (country). Such systems describe non-adiabatic systems that are a real phenomenon of the global system. On the other hand, it can be easily seen that the availability of purchasing power of $A$ to $B$ decreases continuously since the rate of $C$ tends to decrease over time under the inflationary condition. This case, for example, can be observed on the $x\simeq t^{h}$, where a decrease in $C$ is observed at the rate $\frac{h}{t}$ over time when $h>1$. Therefore, it is seen that the economic system has a dynamic structure with $x$ scaling.
Figure 1: Representation of any subsystem chosen in the $E$ economic system. Points AA and $B$ in the subsystem are scaled by $x$. The AA unit of density (a buyer) defined in $A$ is represented by kinetic energy (purchasing power). Point $B$ is the potential energy density representing movable-immovable goods or various assets. During inflation, there exists pressure caused by the scalar field towards the outside of the system $E$, and the system always has a dynamic and expanding feature. While it shows continuous growth on the $x$ scale in the expanding field, it should be controlled with fiscal policies suitable for the system for purchasing power and goods price balance.
@@
The energy density and pressure of this homogeny and dynamic system are given as followsR41 ; R42 ,
$\rho=\frac{\dot{\varphi}^{2}}{2}+V(\varphi),\,\,\,\,\,\,\,\,\,\,p=\frac{\dot{\varphi}^{2}}{2}-V(\varphi)$ (1)
Here, the kinetic energy is the purchasing power at the $A$ point, and the $V(\varphi)$ potential energy represents all the movable and immovable goods at the $B$ point. It is clear that the decline in $C$ reduces the accessibility and so defines the relationship between purchasing power and goods (price). In this regard, the quantities $A$ and $B$ are important objects of the system for our discussion. It is useful to note that inflation occurs when $\frac{\dot{\varphi}^{2}}{2}<V(\varphi)$. The economic case is at a desired level (economics equilibrium) when $\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$. The first ($\frac{\dot{\varphi}^{2}}{2}<V(\varphi)$) is when the price of goods is high but purchasing power is low. This represents a phase in which the $x$ scale continuously increases very rapidly. The second one indicates the existence of an economic equilibrium. In the second case, there is a slow increase in the $x$ scale, which is a necessary condition for a developed country’s economy. Therefore, the first case produces negative pressure corresponding to inflation (expanding of the field), and the second one gets positive pressure corresponding to purchasing power (deceleration of the field). On the other hand, we proceed by defining a modulating parameter, $\gamma=\frac{p_{\varphi}}{\rho_{\varphi}}$, which also denotes a state of the equation. While it is an inflation case with $\gamma<0$, the positive definition of $\gamma$ shows a non-inflation case (decelerating case in the between points).
@@ II.1 Adiabatic process
Figure 2: The relationship between the distance scale and time is given. An exponential increase in the $x$ scale corresponds to hyperinflation. In the $\frac{\dot{\varphi}^{2}}{2}\ll V(\varphi)$ expression, purchasing power (kinetic energy) becomes almost negligible, and the price increase is caused by high potential (movable and immovable property, various assets).
The high potential indicates the high price of material resources that have not yet been converted into monetary (kinetic energy) resources. An exponential increase in the $x$ distance scale indicates a decrease in purchasing power. This unstable situation can be eliminated to some extent by reducing the demand for the purchase of goods, which leads to a decay of the potential, that is, to a decrease in price. However, in such an adiabatic system (that is, there is no monetary input and output to the system), the potential energy (price) can be reduced by increasing production, or, as we emphasized, the price of resources can be reduced by decreasing demand. An important point to be emphasized here is that forcing the system into an imaginary non-adiabatic state is not a solution. Because, even if a heat transfer (monetary resources that can come from outside the country, for example, transferring money to the system with interest) increases the kinetic energy of the system in the short run, it will lead to heavier inflation in the long run. This can be seen in Figure 2 from the peaking of the graph with increasing time. In addition, non-real temporary monetary resource inflow for the system forces the inflation field to a heterogeneous structure in real terms. In the long run, increasing the virtual kinetic energy in the system will lead to the draining of the resources of the system with interest and lead to more imbalance in the densities within the area. Because monetary resources will lead to the gain of one side and more exploitation of the other side, which will result in the emergence of an inhomogeneous area. In real terms, the purchasing power of density units (persons) whose kinetic energy decreases more will weaken and high inflation will continue. These are some conditions that apply to the adiabatic system. Unit densities in adiabatic systems appear to be a more realistic way to reduce demand (making economic savings) for goods, reduce inflation, and bring the economy back to reasonable equilibrium. In this case, the seller will have to reduce the price to convert his movable-immovable or various assets into money, which is one of the most important factors in reducing inflation. On the other hand, by revealing the natural resources in the system, by increasing the resources, the price can be lowered and economic stability can be achieved. Natural resources will lead to the degradation of potential and an increase in monetary (kinetic energy) valuation. Indecision will continue, as wasting resources is contrary to the approach and policy of buying fewer goods. It should be noted here that being satisfied with the goods in hand will either be implemented by the units themselves or economic policies will force the system to do so. The change in the rate of inflation in this phase is equal to $C=\rho_{0}$. If we fix the value of $\rho_{0}$ at a very large value, inflation becomes continuous with an approximately unchanged value of $C$. Therefore, within the framework of reasonable policies, this rate of change parameter should take a time-dependentR42 form that will undergo time evolution. In this direction, in developed economic systems, the system should reach the equilibrium point after inflation. This equilibrium point can be thought to be possible if the purchasing power is equal to the potential, $\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$, corresponding to the point $\gamma\simeq 0$. But, the solution of the eq. (10) around at this point, $\gamma\simeq 0$, is an empty set. Therefore, although $K=V$ for adiabatic systems, as a result of reasonable policies (example of reducing the demand for goods by saving), we obtain a linear solution as follows at the point $\gamma\simeq-\frac{1}{3}$ with the falling pressure of the system.
@@
with real part ($C_{1}=0$), we have $x\sim tC_{2}$. The ratio parameter is in the form $C=\frac{1}{t}$, which exhibits a decreasing behavior with time. As a result, fiscal policies that would require a long time under the constraint of the adiabatic system (economy or fiscal restraint policies) may result in lowering inflation. If $t$ is taken as years, it may take more than 10 years for inflation to be expressed in single digits. As we will see in the next chapter, a reasonable economic equilibrium point may emerge at $Kinetic(K)=Potential(V)$, and a decrease in inflation can be observed in less than 10 years with inter-system trade, except for fiscal constraints or the position of unit densities to settle for less goods.
@@ II.2 Non-adiabatic process
$3\ddot{x}=4\pi\rho_{0}(1+3\gamma)x^{-3(1+\gamma)}.$ (15)
@@
where we set $C_{1}=C_{2}=0$, so, an increase in $x$ is parabolic, indicating very low economic purchasing power.
@@
The proportional parameter of this type of inflationary field behaves as $C=\frac{2}{t}$. Even if the inflation parameter $C$, which starts in a zero time instance, returns to its natural appearance after 20 years, it reveals a deep deformation situation that ruins the field due to the size of the $x$ scale. Therefore, in such an inflationary field, it is necessary to intervene in the system with reasonable fiscal policies and to pass the inflationary phase to a new phase. In this way, a stable course can be followed by keeping the economy under control without any deformation. Therefore, when we apply the $K=V$ or $w\simeq 0$ approximation that we discussed for economic equilibrium to the equation, we get the following equation.
$x=\pm(\frac{-4\pi\rho_{0}+3t^{2}C_{1}^{2}C_{2}+6tC_{1}^{2}C_{2}+3C_{1}^{2}C_{2}^{2}}{3C_{1}})^{\frac{1}{2}}.$ (17)
@@
Bolivia 1970-1985
Peru 1970-1990
Israel 1970-1985
Brazil 1970-1994
Nicaragua 1985-1991
Hungary 1946 Jan-1946 Jul
Germany 1922 Nov-1923 Oct
@@
With the decrease in demand for the density (energy, that is, money) units that describe the system, the kinetic energy, that is, the purchasing power will increase over time. Another possible situation is the realization of the political intervention in the system to reduce the demand for the purchase of goods. Since this policy will reduce purchasing demand over time, the consumer’s decreasing list of needs will increase kinetic energy over time, thus increasing purchasing power. However, in the non-adiabatic process, external monetary investments in the system or the existence of rich resources (potential assets that can be converted into money, that is, kinetic energy) in the field of the system will push the process to an economic balance over time, which our solutions in this topic have shown. Table 1 shows some countries experiencing hyperinflationary periods. When considered yearly, it is observed that there exists an inflation process of about 10 years. This periodic process is in harmony with our calculations. On the other hand, under the condition of applying reasonable financial policies, the economic system in Turkey for example, which started in 2017 and evolved into a hyperinflationary process, may be expected to complete its natural course after 7 years and decrease to a single-digit after 11 years (around 2028 date).
@@ III Closing remarks
There exist many symmetries in the universe (e.g. spatial and temporal translation symmetries, parity, and charge symmetries, etc.). Especially in physical science, symetry describes that when any observable quantity is considered, some properties of the system do not change under certain transformations (for example, Noether symmetries). In other words, symmetry tells us that the internal property of a physical system related to physics or mathematics does not change with the change of some factors. In social science, which is also a part of the universe, evaluating the welfare of society, the interaction of the units that make up the social system (including political interactions) and its development in terms of its compatibility with the laws of theoretical physics, can offer a realistic perspective. On the other hand, the physical development of a plant or human, the growth of the circle of ideas, or the expansion of the intellectual sphere of influence (just as the expansion of the universe) refer to observable symmetries. In this study, which we handled mathematically according to this point of view, which we call ”developmental symmetry” in the universe, we discuss inflation as an ever-expanding system (similar to the dynamic structure in developed economies) with its scalar field representation. In this direction, we investigate the economics of a system from an inflationary perspective within the framework of the Raychaudhuri equation. We have obtained a continuity equation based on the first law of thermodynamics. For the adiabatic and non-adiabatic processes of this equation, we obtain some solutions corresponding to the inflationary case for the system. As can be seen in Figure 1, the main point we consider in the study is the purchasing power, which is described by point $A$, and the purchased goods, which is identified by point $B$. It should be noted that in a developed economic system under consideration, the $x$-scale can never reach a value of zero. Finding an $x$ value in developed economies is important for the continuity and development of the economic system. Therefore, it is inevitable for a dynamic system (developed economic system) to have a constant value ($\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$) of purchasing power and price of goods for an exchange of approximately $x$ value to take place. An increase in the $x$ scale means that the availability of the good decreases, that is, the purchasing power ($\frac{\dot{\phi}^{2}}{2}$) decreases or the price of the good $V(\phi)$ is high. We represented this situation with $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$, therefore, solutions in which hyperinflation occurs in the case of $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$. As emphasized, the equilibrium point is the case of $\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$, which corresponds to a deflationary area that reduces the rate of hyperinflation. Since there is no energy (monetary) flow with external systems for an adiabatic economic system, it has been revealed that it can get rid of inflation in more than 10 years after it evolves into an inflation area of $\gamma=-\frac{1}{3}$ with reasonable financial policies within the system. Also considering the duration of reasonable policies, it will probably take more than 20 years to get out of the inflationary field. For the more realistic non-adiabatic process (the system with energy-money flow), we obtained mathematical results that are quite compatible with the inflation processes seen in Table 1.
@@
On the other hand, the main reason why some countries in Table 1 left the hyperinflation area in less than 11 years can be shown as the inflow of money into the country. For example, the cash flow from the international monetary fund to the country arising from production, natural resources, technological infrastructure and the sound economic policies created as a result can be shown. The reasons for the inflation that continues for more than 11 years on average may be wrong fiscal policies and management of the system (country) as an adiabatic system. In our current study, the 11-year average period of the inflation process is obtained within the framework of ”developmental symmetry” for non-adiabatic systems. After an average of 11 years of hyperinflation, the economic system is evolving into a stable state (single-digit percentages). As we have emphasized, one of the most effective ways to stop hyperinflation in such an economic system created identically (symmetrically) to mathematical tools is to ensure production in the system. Whether such a system is adiabatic or not (whether there is energy-money flow or not) may no longer be important. The use of natural resources with the right policies or the production of technological parameters that may meet the needs can lead to a reasonable inflation rate without the flow of money with interest to the inflation area where purchasing power drops to almost zero. In another separate study, it can be discussed whether inflation can be cut or not by finding the creation of the scalar field density in the field. Because density is a quantity that includes kinetics and potential due to $E\sim\rho$. Units of density created in the field by the political ground controlling the field represents the unity of purchasing power (kinetic energy, point $A$ in Figure 1) and movable immovable property (potential energy, point $B$ in Figure 1). Therefore, it is clear that this unity emerging everywhere in the field will have a decreasing effect on inflation.

### v2

@@ I Introduction
In this study, it is shown that Raychaudhuri equation R32 ; R33 ; R34 ; R35 can describe an economic system without going into deep mathematical proofs. The solutions of the equation for adiabatic and non-adiabatic processes that occur in the energy conservation equation, which are considered in the thermodynamic framework, are investigated. At this point, it has been observed that the adiabatic process in which hyperinflation emerged can be eliminated by certain economic processes and monetary policies directing the system. In this direction, we consider the economic system as a field that can be defined as a scalar field $\phi$. The existence of a minimum inflation in this field is a necessity even at the beginning. The Raychaudhuri equation appears as an equation describing such an economic system that can bend and expand continuously. When we say bending or curvature of the field, we mean whether the unit (point A in Figure 1) with kinetic energy (monetary purchasing power) can reach a point with potential energy (movable-real estate, valuables) within the field. In this context, the present study aims to formulate an analytical framework in which macroeconomic behavior can be interpreted through the geometrical properties of the economic field. This approach provides a novel perspective for understanding inflationary dynamics and market evolution in terms of curvature and expansion processes. Thus, the proposed analogy offers a new interpretive framework connecting economic stability and inflationary behavior with field dynamics, suggesting that economic equilibria may be viewed as geometrical configurations within this scalar field. Therefore, this analogy provides a foundation for modeling economic evolution as a geometric process, allowing inflationary behavior and market expansion to be understood within a unified thermodynamic and relativistic framework.
@@ II The Raychaudhuri Equation and Inflation
In general, inflation is a concept used to explain the continuous increase in the average price level within an economy. At the same time, this increase implies a decline in the value of money (Reserve Bank of New Zealand R36 ). In the mathematical constructs of economics, rather than relying solely on theoretical proofs, conceptual economic phenomena are structured within the framework of general system theory.
It should be noted that the present work does not aim to establish a strict physical equivalence between cosmological and economic variables. Instead, the analogy developed here is used phenomenologically to describe the qualitative behavior of economic expansion and contraction processes through the lens of dynamic systems.
With these intuitive and reasonable assumptions, economic systems are examined as evolving entities that can be analyzed using methods inspired by physical sciences. We will mathematically restructure the dynamics of inflation, defined as the persistent rise in the prices of movable and immovable goods R31 , not as a process limited to a subsystem, but as an intrinsic property of the general structure of the economy, represented within a continuous field-like framework.
Here, the economic variable $x(t)$ represents the relative distance between purchasing power and the price level, analogous to a system-wide growth parameter describing overall economic expansion. This definition allows inflation to be expressed in terms of the time evolution of this field.
In this framework, the purchasing power is represented by point $A$ in the field, and the price of goods by point $B$. The distance between both points is scaled by $x$. The availability of $B$ in the economic field (i.e., the presence of goods and services) is assumed to prevent deflationary collapse. The system is considered to be in an inflationary phase when the acceleration term satisfies $\ddot{x}>0$, and in a deflationary phase when $\ddot{x}<0$. An increase in the $x$ scale under the condition $\ddot{x}>0$ indicates a decrease in purchasing power, thus characterizing inflationary behavior.
Consequently, the acceleration of the economic field can be viewed as a measure of inflationary pressure, where fiscal and monetary interventions act as external forces capable of modulating the field’s expansion rate.
In the literature, the relative rate of change in the price of goods is generally defined as
$C=\frac{\dot{p}}{p}$
R37 ; R38 ; R39 ; R40 . According to the present framework, an alternative definition of the inflation rate can be expressed as
$C=\frac{\dot{x}}{x}$
which shows how an economic system or equivalently, an inflation field evolves over time. In such a system, the establishment of a developed economic equilibrium depends on the system’s ability to convert production value, represented by $B$, into purchasing power at point $A$.
In an inflationary context, the conversion of movable or immovable assets (including technological goods) into monetary resources often occurs through export activities or external resource inflows. These represent non-adiabatic processes within the global economic environment. On the other hand, it can be observed that the accessibility of $A$ to $B$ decreases continuously, since $C$ tends to decline over time under inflationary conditions. This behavior can, for instance, be modeled as
$x\simeq t^{h}$
where a decrease in $C$ occurs at a rate
$\frac{h}{t}$
when $h>1$. Therefore, the economic system exhibits a dynamic structure characterized by the scaling behavior of $x$.
Figure 1: Conceptual representation of a subsystem within the economic system $E$. Points AA and $B$ indicate two representative positions in the system, scaled by $x$. The element at AA corresponds to purchasing power, while $B$ represents the value of goods or assets. During inflationary expansion, systemic forces act outward, reflecting a general increase in the scale of economic transactions. Fiscal and monetary interventions function as regulatory mechanisms that help maintain balance between purchasing power and asset prices within this dynamic structure.
@@
The energy density and pressure of this homogeneous and dynamic economic field are given as follows R41 ; R42 ,
$\rho=\frac{\dot{\varphi}^{2}}{2}+V(\varphi),\,\,\,\,\,\,\,\,\,\,p=\frac{\dot{\varphi}^{2}}{2}-V(\varphi)$
Here, the kinetic term $\frac{\dot{\varphi}^{2}}{2}$ corresponds to the purchasing power at point $A$, while the potential energy $V(\varphi)$ represents the accumulated value of movable and immovable goods at point $B$. A decrease in $C$ reflects reduced accessibility between purchasing power and goods prices. Inflation occurs when
$\frac{\dot{\varphi}^{2}}{2}<V(\varphi)$
whereas an economic equilibrium is achieved when
$\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$
The former corresponds to a phase of rapid expansion in the $x$-scale, characterized by high prices and low purchasing power, while the latter represents a stable, slowly expanding economy typical of developed systems. Therefore, the first case produces negative pressure associated with inflation (field expansion), whereas the second results in positive pressure reflecting purchasing power and economic stabilization.
A state parameter can be defined as
$\gamma=\frac{p_{\varphi}}{\rho_{\varphi}}$
which denotes the general state of the system. When $\gamma<0$, the system is in an inflationary (expanding) phase, while a positive $\gamma$ indicates a non-inflationary or decelerating condition.
Therefore, the analytical framework developed here allows inflation to be interpreted as a curvature-like expansion within an abstract economic field. This analogy highlights the role of fiscal and monetary interventions as control parameters that regulate the dynamic balance between purchasing power and asset values. While the proposed model remains phenomenological, it provides a conceptual bridge connecting macroeconomic behavior with the systemic dynamics of complex evolving structures observed across both physical and social domains.
@@ II.1 Adiabatic process
Figure 2: The relationship between the scale variable $x$ and time $t$. The exponential increase in $x$ represents the hyperinflationary phase of the system. In this regime, purchasing power, interpreted as the system’s kinetic component, becomes almost negligible, while the rapid increase in prices is driven by accumulated asset values and capital concentration, corresponding to the potential component of the system.
The high potential indicates high prices of material or financial resources that have not yet been converted into monetary resources, that is, into the system’s kinetic component. An exponential increase in the scale variable $x$ indicates a loss of purchasing power. This unstable situation can be mitigated by reducing demand for goods, which in turn leads to a decline in potential, that is, a decrease in prices.
In an adiabatic system meaning no external monetary inflow or outflow the potential (price level) can be reduced either by increasing production or by lowering demand. It should be emphasized that artificially forcing the system into a non-adiabatic state is not an effective solution. Although a temporary monetary inflow (for example, capital entering the system through interest-based borrowing) can increase the kinetic component (purchasing power) in the short term, it ultimately leads to stronger inflation in the long run. This is reflected in Figure 2, where the curve peaks over time.
Furthermore, temporary external monetary inputs distort the homogeneity of the system by introducing artificial kinetic energy, which eventually drains resources through interest payments and creates structural imbalance. Monetary inflows enrich one segment while exploiting another, leading to an inhomogeneous field where purchasing power continues to erode. Hence, sustained inflation persists, especially for units whose kinetic energy (purchasing capacity) declines more rapidly.
Under adiabatic conditions, the most realistic path toward restoring equilibrium involves reducing demand encouraging economic savings and consumption restraint. This drives sellers to lower prices to convert their assets into money, which is a key mechanism for reducing inflation. In addition, increasing the availability of natural or productive resources decreases potential (prices) and enhances monetary valuation, moving the system toward stability.
However, wasting or underutilizing resources contradicts the policy direction of minimizing consumption. Therefore, satisfaction with existing goods either voluntarily or as a result of policy pressure becomes a natural adjustment mechanism.
The rate of inflation in this regime is given by $C=\rho_{0}$. If $\rho_{0}$ is fixed at a large value, inflation remains nearly constant over time. Thus, under rational policy frameworks, the rate of change $C$ should take a time-dependent form R42 to allow for gradual adjustment. In developed economies, the system tends to reach an equilibrium where purchasing power equals potential, $\frac{\dot{\varphi}^{2}}{2}\simeq V(\varphi)$, corresponding to $\gamma\simeq 0$. However, the solution of Eq. (10) near this point ($\gamma\simeq 0$) is an empty set.
Therefore, although $K=V$ holds approximately in adiabatic systems, under rational fiscal policies–such as demand reduction through savings–the system evolves toward a linear behavior at $\gamma\simeq-\frac{1}{3}$, where pressure gradually falls:
@@
Taking the real part ($C_{1}=0$), we obtain $x\sim tC_{2}$. The ratio parameter takes the form $C=\frac{1}{t}$, showing a decreasing trend over time. As a result, fiscal restraint policies that are maintained over long periods can effectively lower inflation.
If $t$ is measured in years, this model suggests that inflation could decline to single-digit levels over a period longer than ten years. As will be discussed in the following section, when moderate inter-system trade is allowed, an economic equilibrium point may be reached at $K=V$, and inflation may decline more rapidly within less than a decade even without strict fiscal constraints.
@@ II.2 Non-adiabatic process
$3\ddot{x}=4\pi\rho_{0}(1+3\gamma)x^{-3(1+\gamma)},$ (14)
@@
where we set $C_{1}=C_{2}=0$, so an increase in $x$ is parabolic, indicating very low economic purchasing power.
@@
The proportional parameter of this type of inflationary field behaves as $C=\frac{2}{t}$. Even if the inflation parameter $C$, which starts in a zero time instance, returns to its natural appearance after 20 years, it reveals a deep deformation situation that ruins the field due to the size of the $x$ scale. Therefore, in such an inflationary field, it is necessary to intervene in the system with reasonable fiscal policies and to pass the inflationary phase to a new phase. In this way, a stable course can be followed by keeping the economy under control without any deformation.
Therefore, when we apply the $K=V$ or $w\simeq 0$ approximation that we discussed for economic equilibrium to the equation, we get the following equation.
$x=\pm\left(\frac{-4\pi\rho_{0}+3t^{2}C_{1}^{2}C_{2}+6tC_{1}^{2}C_{2}+3C_{1}^{2}C_{2}^{2}}{3C_{1}}\right)^{\frac{1}{2}}.$ (16)
@@
Bolivia 1970–1985
Peru 1970–1990
Israel 1970–1985
Brazil 1970–1994
Nicaragua 1985–1991
Hungary 1946 Jan–1946 Jul
Germany 1922 Nov–1923 Oct
@@
With the decrease in demand for the density units (energy, i.e., money) that describe the system, the kinetic energy or purchasing power will increase over time. Another possible scenario is political intervention aimed at reducing demand for the purchase of goods. Such policies reduce purchasing demand gradually, leading consumers to limit their needs, which increases kinetic energy over time and thus enhances purchasing power.
However, in the non-adiabatic process, external monetary investments or the presence of rich resources (potential assets convertible into money, i.e., kinetic energy) within the system will steer the economy toward balance over time, as our solutions have demonstrated.
Table 1 shows some countries that experienced hyperinflationary periods. Considering these intervals, it is observed that inflation typically persists for about 10 years, consistent with our theoretical calculations.
For example, under reasonable financial policies, the economic system in Turkey, which entered a hyperinflationary phase starting in 2017, may be expected to complete its natural course after 7 years and reduce inflation to single digits after approximately 11 years (around 2028).
@@ III Closing remarks
There exist many symmetries in the universe (e.g. spatial and temporal translation symmetries, parity, and charge symmetries, etc.). Especially in physical science, symetry describes that when any observable quantity is considered, some properties of the system do not change under certain transformations (for example, Noether symmetries). In other words, symmetry tells us that the internal property of a physical system related to physics or mathematics does not change with the change of some factors. In social science, which is also a part of the universe, evaluating the welfare of society, the interaction of the units that make up the social system (including political interactions) and its development in terms of its compatibility with the laws of theoretical physics, can offer a realistic perspective. On the other hand, the physical development of a plant or human, the growth of the circle of ideas, or the expansion of the intellectual sphere of influence (just as the expansion of the universe) refer to observable symmetries. In this study, which we handled mathematically according to this point of view, which we call ”developmental symmetry” in the universe, we discuss inflation as an ever-expanding system (similar to the dynamic structure in developed economies) with its scalar field representation. In this direction, we investigate the economics of a system from an inflationary perspective within the framework of the Raychaudhuri equation. We have obtained a continuity equation based on the first law of thermodynamics. For the adiabatic and non-adiabatic processes of this equation, we obtain some solutions corresponding to the inflationary case for the system. As can be seen in Figure 1, the main point we consider in the study is the purchasing power, which is described by point $A$, and the purchased goods, which is identified by point $B$. It should be noted that in a developed economic system under consideration, the $x$-scale can never reach a value of zero. Finding an $x$ value in developed economies is important for the continuity and development of the economic system. Therefore, it is inevitable for a dynamic system (developed economic system) to have a constant value ($\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$) of purchasing power and price of goods for an exchange of approximately $x$ value to take place. An increase in the $x$ scale means that the availability of the good decreases, that is, the purchasing power ($\frac{\dot{\phi}^{2}}{2}$) decreases or the price of the good $V(\phi)$ is high. We represented this situation with $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$, therefore, solutions in which hyperinflation occurs in the case of $\frac{\dot{\phi}^{2}}{2}\ll V(\phi)$. As emphasized, the equilibrium point is the case of $\frac{\dot{\phi}^{2}}{2}\simeq V(\phi)$, which corresponds to a deflationary area that reduces the rate of hyperinflation. Since there is no energy (monetary) flow with external systems for an adiabatic economic system, it has been revealed that it can get rid of inflation in more than 10 years after it evolves into an inflation area of $\gamma=-\frac{1}{3}$ with reasonable financial policies within the system. Also considering the duration of reasonable policies, it will probably take more than 20 years to get out of the inflationary field. For the more realistic non-adiabatic process (the system with energy-money flow), we obtained mathematical results that are quite compatible with the inflation processes seen in Table 1. This analogy between thermodynamic balance and economic equilibrium also reflects the collective behavioral tendencies of economic agents–such as consumption habits, trust in fiscal policy, and market expectations–that determine how energy (money) circulates within a society.
@@
On the other hand, the main reason why some countries in Table 1 left the hyperinflation area in less than 11 years can be shown as the inflow of money into the country. For example, the cash flow from the international monetary fund to the country arising from production, natural resources, technological infrastructure and the sound economic policies created as a result can be shown. The reasons for the inflation that continues for more than 11 years on average may be wrong fiscal policies and management of the system (country) as an adiabatic system. In our current study, the 11-year average period of the inflation process is obtained within the framework of ”developmental symmetry” for non-adiabatic systems. After an average of 11 years of hyperinflation, the economic system is evolving into a stable state (single-digit percentages). As we have emphasized, one of the most effective ways to stop hyperinflation in such an economic system created identically (symmetrically) to mathematical tools is to ensure production in the system. Whether such a system is adiabatic or not (whether there is energy-money flow or not) may no longer be important. The use of natural resources with the right policies or the production of technological parameters that may meet the needs can lead to a reasonable inflation rate without the flow of money with interest to the inflation area where purchasing power drops to almost zero. In another separate study, it can be discussed whether inflation can be cut or not by finding the creation of the scalar field density in the field. Because density is a quantity that includes kinetics and potential due to $E\sim\rho$. Units of density created in the field by the political ground controlling the field represents the unity of purchasing power (kinetic energy, point $A$ in Figure 1) and movable immovable property (potential energy, point $B$ in Figure 1). Therefore, it is clear that this unity emerging everywhere in the field will have a decreasing effect on inflation. This highlights that maintaining the balance between purchasing power and asset value throughout the economic field is crucial for mitigating hyperinflation and ensuring sustainable economic growth.